OpenStudy (anonymous):

PLEASE HELP ME What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = -83? -200 -197 -194 -191

OpenStudy (cwrw238):

you first need to find the common difference d then 35th term = a1 + ( 35 - 1)*d

OpenStudy (cwrw238):

you can find d by calculating (a17 - a1) / x what do you think x has to be?

OpenStudy (anonymous):

i just need the answer, im taking a timed quiz

OpenStudy (anonymous):

i dont know how to solve that

OpenStudy (cwrw238):

you can calculate a17 - a1 = -83 - 13

OpenStudy (cwrw238):

and x = 17 - 1 = 16 terms

OpenStudy (anonymous):

soo? after that

OpenStudy (anonymous):

-83 = 13 + 16d ...and.. -96 = 16d d = -6 So... a_35 = 34(-6) + 13 = -204 + 13 = -191

OpenStudy (anonymous):

HURRRY MY TEST IS ALMOST OVER

OpenStudy (anonymous):

OKAY THANK YOU! :) SO MUCH.

OpenStudy (anonymous):

ALSO What is the sum of a 17-term arithmetic sequence where the first term is 6 and the last term is -90?